Two charrged sphere seprated in air at a distance d exert a force F on each other If they are immersed in a liquid of dielectric constant 2 then what is the force

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

for air K=1

when spehere are placed in air F= 1/4∏ε  q1 *q2 /d2

when sphere are placed in dielectric (k=2)

F1 = 1/4∏εK  q1*q2/d2

      = 1/k [F]

       =F/k

    = F/2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin